OpenStudy (king):

By what number less than 2 must be multiplied so that the last 3 figures on the right end of the product may be 791

OpenStudy (king):

igs hellp man....atleast try

OpenStudy (king):

@JamesJ ,@Tomas.A ,@Diyadiya help guys help!!!

OpenStudy (king):

help callisto!

OpenStudy (king):

this frm real numbers chapter

OpenStudy (lgbasallote):

oh man...im weak with these things @_@

OpenStudy (king):

@Callisto u are the only one with a chance of doin this...so do u know how to do it?

OpenStudy (callisto):

Sorry i don't even understand the question :( what number less than 2 <-- does it mean a single number?

OpenStudy (king):

yeah there is some number...its related to HCF AND LCM THROUGH EUCLID'S DIVISION LEMMA

OpenStudy (callisto):

Sorry haven't learnt that :(

OpenStudy (anonymous):

Word it right, you worded it so complex.

OpenStudy (king):

its a problem in my phase-4 fiitjee package..i gave it to u the way it is worded

OpenStudy (king):

but i dont understand answer is 871

OpenStudy (king):

oh sry wait its by what number less than 2must be multiplied with 43521.....

OpenStudy (king):

i by mistake frgt to mention 43521....:/

OpenStudy (anonymous):

Oh

OpenStudy (king):

so?

OpenStudy (anonymous):

Let x be the number, (435*100,000 + 521)*x = adasd...fsh791 So, last digit of x must be 1. The answer is clearly 871, but I did this problem with logic and it's really tough to explain logic unless your communication skills are really impressive. So... I am gonna tag @FoolForMath and @Mr.Math to see if they know some trick or the method you mentioned above.

OpenStudy (anonymous):

btw basic multiplication is what I used.

OpenStudy (king):

its 43521

OpenStudy (king):

not 43500521

OpenStudy (anonymous):

(435*1000 + 521)*x = adasd...fsh791

OpenStudy (anonymous):

My bad

OpenStudy (king):

its 43*1000 +521 or 435*100+21 rite?

OpenStudy (king):

435*1000 + 521)=435521

OpenStudy (anonymous):

yeah, again my bad :/

OpenStudy (king):

oh and u can use any method......

OpenStudy (anonymous):

but it doesn't matter, just get the 521 separately and then multiply it my x (the assumed number).